OpenStudy (anonymous):

find all angles θbetween 0° and 2pi satisfying the given equation sqrt2 /2

OpenStudy (anonymous):

What is the equation?

OpenStudy (anonymous):

Use the figure to find all angles between and satisfying the given condition (Use symbolic notation and fractions where needed. Give your answer in the form of comma separated list.)

OpenStudy (anonymous):

between 0 to 2pi of the given condition sin = sqrt2 /2

OpenStudy (anonymous):

Oh so the equation is: \[sin(\theta)=\frac{\sqrt{2}}{2}\] And they need you to find the angle \(\theta\) where \(\{0\leq\theta\leq2\pi\}\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Okay. So do you remember the CAST rule? it states where each of the ratios will be positive. It looks like this:  So the S,T, and C represent the sin, cos, and tan ratios (respectively). The A represents "all". From this we can gather that sin will be positive in the First quadrant (top right section) and the second quadrant (top left section). Make sense so far?